Solve hte indicated system of equations by graphing. Interpert your graph and indicate the solution.
3x - y = 9
2x + y = 11
For part of it I have: 5x = 20 x = 4

Answer by fractalier(6550)

You are okay with
3x - y = 9
2x + y = 11
For part of it I have: 5x = 20 x = 4
Now plug this value of x into either equation to get y...
3(4) - y = 9
12 - y = 9
y = 3
So the solution is (4, 3)